3x3 - 36x2 + 96x
1
Expert's answer
2012-10-25T09:15:34-0400
we can factorize it
3x3 - 36x2 + 96x=3*x*(x^2-12x+32)=3x(x-4)(x-8)
so
roots are 0, 4, 8